# Convergence and long-time behavior of finite volumes for a generalized Poisson-Nernst-Planck system with cross-diffusion and size exclusion

### v1

@@ 5. Long-time asymptotic of the scheme
for $\sigma=K|L\in\mathcal{E}_{\text{int}}$, where we have set
@@
We are moreover interested in the convergence of $(U^{n}_{\mathcal{T}},\phi_{\mathcal{T}}^{n})$ towards $(U^{\infty}_{\mathcal{T}},\phi_{\mathcal{T}}^{\infty})$ as $n$ goes to +$\infty$+ ∞.
@@
[…]
such that,
@@
Owing to (42) and the non-negativity of the discrete dissipations $\mathcal{D}^{n}_{\mathcal{T}}$, the sequence ${(\mathcal{H}_{\mathcal{T}}^{n})}_{n\geq 0}$ is decreasing, and moreover bounded thanks to Lemma 3.6. Therefore, it converges towards some finite limit $\mathcal{H}_{\mathcal{T}}^{\infty}$, while $\mathcal{D}_{\mathcal{T}}^{n}$ has to tend to 00 as $n$ goes to +$\infty$+ ∞. Then so does $\mathcal{D}_{i,\sigma}^{n}$, cf. (43), for all $i=1,\dots,I$ and all $\sigma=K|L\in\mathcal{E}_{\text{int}}$. Since $\left(U_{\mathcal{T}}^{n},\phi_{\mathcal{T}}^{n}\right)_{n\geq 0}$ is included in a bounded (thus relatively compact) set, the LaSalle invariance principle yields the existence of (at least) an accumulation point $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$, the energy of which being equal to $\mathcal{H}_{\mathcal{T}}^{\infty}$. Then it follows from the discrete energy / energy dissipation inequality (42) that the dissipation $\mathcal{D}_{\mathcal{T}}^{\infty}$ corresponding to $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$ is equal to 00, then so are all the edge dissipation contribution $\mathcal{D}_{i,\sigma}^{\infty}$ which are defined by (43) but with $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$ instead of $\left(U_{\mathcal{T}}^{n},\phi_{\mathcal{T}}^{n}\right)$. Since $\mathcal{D}_{i,\sigma}^{\infty}=0$ iff the corresponding flux $F_{i,K\sigma}^{\infty}$ defined by (80) vanishes for all $i=1,\dots,I$ and $\sigma=K|L\in\mathcal{E}_{\text{int}}$, one gets that $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$ is a steady solution, and that the convergence (84) holds true.
@@
This can be seen using the two formulations of the fluxes (80) and (31) at the limit $n\to\infty$. The formula (80) clearly yields (i) to (iii) if any of the involved volume fraction vanish. If they are all positive then (31) is well-defined at the limit $n\to\infty$ and yields (iv).
Assume for contradiction that there exists $K\in\mathcal{T}$ such that $u_{0,K}^{\infty}=0$, then $F_{i,K\sigma}^{\infty}=0$ for all $i$ requires that either $u_{i,K}^{\infty}=0$ for all $1\leq i\leq I$, which is incompatible with (81), or that $u_{0,L}^{\infty}=0$ for all neighboring cell $L$ sharing an edge with $K$. An induction shows relying on the connected character of $\Omega$ shows that $u_{0,\mathcal{T}}^{\infty}=0$, which contradicts (82). The cases (i) and (iii) are therefore impossible, and $u_{0,\mathcal{T}}^{\infty}>0$.
Assume now that $u_{i,K}^{\infty}=0$ for some $K\in\mathcal{T}$, then as $u_{0,K}^{\infty}>0$, a similar reasoning shows that $u_{i,L}^{\infty}$ also has to be equal to 00 for all neighboring cell of $K$. Then $u_{i,\mathcal{T}}^{\infty}=0$, contradicting again (82). Therefore, alternative (ii) is not feasible, and (iv) has to hold true. This concludes the proof of Proposition 5.1. ∎
@@
[…]
with $v_{i}$ defined in (18). Incorporating this relation in (78) provides the following discrete counterpart to the modified Poisson-Boltzmann equation (20):
@@
On readily checks that
@@
(87) $A$ vanishes if and only if ${y}_{\mathcal{T}}={\tilde{y}}_{\mathcal{T}}$.
@@
the term $B$ rewrites as
(88) $B=\sum_{K\in\mathcal{T}}m_{K}\left\langle\frac{\partial H^{*}}{\partial{r}_{i}}({R}_{K})-\frac{\partial H^{*}}{\partial{r}_{i}}({\tilde{R}}_{K}),{R}_{K}-{\tilde{R}}_{K}\right\rangle.$
@@ 6. Numerical results
The nonlinear system corresponding to the scheme is solved thanks to a Newton-Raphson method with stopping criterion ${\|\mathcal{F}_{\mathcal{T}}^{n}((U_{K}^{n})_{K\in\mathcal{T}},(\phi_{K}^{n})_{K\in\mathcal{T}})\|}_{\infty}\leq 10^{-10}$, the components of $\mathcal{F}_{\mathcal{T}}^{n}$ being given by the left-hand side of (25).
@@ 6.1. Convergence under grid refinement
A reference solution is computed on a grid made of $1638400$ cells and with a constant time step $\tau=10^{-3}$, to which are compared solutions computed on successively refined grids but with the same constant time step. The profile of the solution at the final time $T=1$ is depicted on Figures 2. The relative space-time $L^{1}$ error is plotted as a function of the number of cells on Figure 3, showing some second order accuracy in space, as specified in the introductory discussion. The number of Newton iterations required to solve a time step remains between 6 for the very first iterations and 2 for larger times is mainly insensitive to the mesh size.
@@ 6.2. Long-time behavior of the scheme
We now focus on the long-time behavior of the scheme, and more specifically on the convergence towards its unique steady state. Our study relies on a two-dimensional test case, with $I=3$ charged species (plus the solvent $u_{0}$). The corresponding charges are $z_{1}=2$, $z_{2}=1$, $z_{3}=-1$ (and $z_{0}=0$ as the solvent is electrically neutral in our model), while the diffusion coefficients are set to $D_{1}=1$, $D_{2}=2$ and $D_{3}=2$.
The geometry is the following : $\Omega$ is the unit square $(0,1)^{2}$, the boundary of which being split into $\Gamma^{D}=\{y=1\}\times\{0\leq x\leq 1/2\}$ and its complement $\Gamma^{N}$. The electric potential is set to 00 on $\Gamma^{D}$, i.e. $\phi^{D}=0$.
@@
We run our scheme with a constant time step $\tau=10^{-4}$ and for two different Debye length until a final time $T=3$, and look for the evolution of the relative energy $\mathcal{H}_{\mathcal{T}}^{n}-\mathcal{H}_{\mathcal{T}}^{\infty}$ along time. The relative energy is decaying for all the curves plotted on Figure 6, but the velocity at which the decay occurs varies strongly depending on the Debye length and on the initial profile. In the electroneutral configurations, we mainly observe that higher quantities of $u_{0}$ yield a faster convergence. This is expected as the dissipation (14) contains a $u_{0}^{2}$ prefactor, in opposition to a mere $u_{0}$ for the classical Poisson-Nernst-Planck system. Then, still in the electroneutral regime, the smaller is $\lambda^{2}$ then the faster is the convergence towards the constant in space equilibrium.
@@
The trend is different for globally charged profiles. For the choice (91) of the initial profile, the convergence becomes extremely slow when $\lambda^{2}$ becomes small. This is due to the fact that, in some areas, the concentration of $u_{0}$ becomes extremely small. The equilibration between the other phases is then extremely slow in these regions. To illustrate this difficulty, we depict on Figure 7 the concentration profiles at $T=3$. The solvent concentration $u_{0}$ is already in good agreement with the equilibrium profile depicted on Figure 5 (see also Figure 8 for a representation of the difference between $U_{\mathcal{T},\tau}(T)$ and $U_{\mathcal{T}}^{\infty}$), but this is not the case of the other concentrations, the evolution of which being stuck in the zones where the concentration of $u_{0}$ is small. Getting a fine quantitative understanding of the long-time behavior of the model goes beyond the scope of this paper and should be investigated in future works.

### v2

@@ 5. Long-time asymptotic of the scheme
for $\sigma=K|L\in\mathcal{E}_{\rm int}$, where we have set
@@
We are moreover interested in the convergence of $(U^{n}_{\mathcal{T}},\phi_{\mathcal{T}}^{n})$ towards $(U^{\infty}_{\mathcal{T}},\phi_{\mathcal{T}}^{\infty})$ as $n$ goes to $\infty$.
@@
(83) $\log\frac{u_{i,K}^{\infty}}{u_{0,K}^{\infty}}+z_{i}\phi_{K}^{\infty}=\mu_{i,\mathcal{T}}^{\infty},\qquad K\in\mathcal{T},\;1\leq i\leq I,$
such that
@@
Owing to (42) and the non-negativity of the discrete dissipations $\mathcal{D}^{n}_{\mathcal{T}}$, the sequence ${(\mathcal{H}_{\mathcal{T}}^{n})}_{n\geq 0}$ is decreasing, and moreover bounded thanks to Lemma 3.7. Therefore, it converges towards some finite limit $\mathcal{H}_{\mathcal{T}}^{\infty}$, while $\mathcal{D}_{\mathcal{T}}^{n}$ has to tend to 0 as $n$ goes to $\infty$. Then so does $\mathcal{D}_{i,\sigma}^{n}$, cf. (43), for all $i=1,\dots,I$ and all $\sigma=K|L\in\mathcal{E}_{\rm int}$. Since $\left(U_{\mathcal{T}}^{n},\phi_{\mathcal{T}}^{n}\right)_{n\geq 0}$ is included in a bounded (thus relatively compact) set, the LaSalle invariance principle yields the existence of (at least) an accumulation point $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$, the energy of which being equal to $\mathcal{H}_{\mathcal{T}}^{\infty}$. Then it follows from the discrete energy / energy dissipation inequality (42) that the dissipation $\mathcal{D}_{\mathcal{T}}^{\infty}$ corresponding to $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$ is equal to 0, then so are all the edge dissipation contributions $\mathcal{D}_{i,\sigma}^{\infty}$ which are defined by (43) but with $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$ instead of $\left(U_{\mathcal{T}}^{n},\phi_{\mathcal{T}}^{n}\right)$. Since $\mathcal{D}_{i,\sigma}^{\infty}=0$ iff the corresponding flux $F_{i,K\sigma}^{\infty}$ defined by (80) vanishes for all $i=1,\dots,I$ and $\sigma=K|L\in\mathcal{E}_{\rm int}$, one gets that $\left(U_{\mathcal{T}}^{\infty},\phi_{\mathcal{T}}^{\infty}\right)$ is a steady solution, and that the convergence (84) holds true.
@@
This can be seen using the two formulations of the fluxes (80) and (31) at the limit $n\to\infty$. The formula (80) clearly yields (i) to (iii) if any of the involved volume fractions vanish. If they are all positive then (31) is well-defined at the limit $n\to\infty$ and yields (iv).
Assume for contradiction that there exists $K\in\mathcal{T}$ such that $u_{0,K}^{\infty}=0$, then $F_{i,K\sigma}^{\infty}=0$ for all $i$ requires that either $u_{i,K}^{\infty}=0$ for all $1\leq i\leq I$, which is incompatible with (81), or that $u_{0,L}^{\infty}=0$ for all neighboring cells $L$ sharing an edge with $K$. An induction shows relying on the connected character of $\Omega$ shows that $u_{0,\mathcal{T}}^{\infty}=0$, which contradicts (82). The cases (i) and (iii) are therefore impossible, and $u_{0,\mathcal{T}}^{\infty}>0$.
Assume now that $u_{i,K}^{\infty}=0$ for some $K\in\mathcal{T}$, then as $u_{0,K}^{\infty}>0$, a similar reasoning shows that $u_{i,L}^{\infty}$ also has to be equal to 0 for all neighboring cell of $K$. Then $u_{i,\mathcal{T}}^{\infty}=0$, contradicting again (82). Therefore, alternative (ii) is not feasible, and (iv) has to hold true. This concludes the proof of Proposition 5.1. ∎
@@
$u_{i,K}^{\infty}=v_{i}(\phi_{K}^{\infty},\boldsymbol{\mu}_{\mathcal{T}}^{\infty}),\qquad K\in\mathcal{T},\;0\leq i\leq I$
with $v_{i}$ defined in (18). Incorporating this relation in (78) provides the following discrete counterpart to the modified Poisson–Boltzmann equation (20):
@@
One readily checks that
@@
(87) AA vanishes if and only if ${y}_{\mathcal{T}}={\tilde{y}}_{\mathcal{T}}$.
@@
the term $B$ is rewritten as
(88) $B=\sum_{K\in\mathcal{T}}m_{K}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{i=1}^{I}}\left\langle\frac{\partial H^{*}}{\partial{r}_{i}}({R}_{K})-\frac{\partial H^{*}}{\partial{r}_{i}}({\tilde{R}}_{K}),{R}_{K}-{\tilde{R}}_{K}\right\rangle.$
@@ 6. Numerical results
The non-linear system corresponding to the scheme is solved thanks to a Newton–Raphson method with stopping criterion ${\|\mathcal{F}_{\mathcal{T}}^{n}((U_{K}^{n})_{K\in\mathcal{T}},(\phi_{K}^{n})_{K\in\mathcal{T}})\|}_{\infty}\leq 10^{-10}$, the components of $\mathcal{F}_{\mathcal{T}}^{n}$ being given by the left-hand side of (25).
@@ 6.1. Convergence under grid refinement
A reference solution is computed on a grid made of $1638400$ cells and with a constant time step $\tau=10^{-3}$, to which are compared solutions computed on successively refined grids but with the same constant time step. The profile of the solution at the final time $T=1$ is depicted on Figures 2. The relative space-time $L^{1}$ error is plotted as a function of the number of cells on Figure 3, showing some second order accuracy in space, as specified in the introductory discussion. The number of Newton iterations required to solve a time step remains between 6 for the very first iterations and 2 for larger times and is mainly insensitive to the mesh size.
@@
Remark 6.1.
An optimal error estimate showing order 2 in space and order 1 in time convergence can be established on cartesian grids if one assumes that there exists a smooth strong solution to the continuous problem, see for instance [12] in a scalar context, or [21] for classical Poisson–Nernst–Planck system. The extension to unstructured grids is then much more delicate, as usual for finite volumes, see for instance [24] for a contribution in this direction. However, the existence of such a strong solution to the problem is an open question.
@@ 6.2. Long-time behavior of the scheme
We now focus on the long-time behavior of the scheme, and more specifically on the convergence towards its unique steady state. Our study relies on a two-dimensional test case, with $I=3$ charged species (plus the solvent $u_{0}$). The corresponding charges are $z_{1}=2$, $z_{2}=1$, $z_{3}=-1$ (and $z_{0}=0$ as the solvent is electrically neutral in our model), while the diffusion coefficients are set to $D_{1}=1$, $D_{2}=2$, and $D_{3}=2$.
The geometry is the following : $\Omega$ is the unit square $(0,1)^{2}$, the boundary of which being split into $\Gamma^{D}=\{y=1\}\times\{0\leq x\leq 1/2\}$ and its complement $\Gamma^{N}$. The electric potential is set to 0 on $\Gamma^{D}$, i.e. $\phi^{D}=0$.
@@
We run our scheme with a constant time step $\tau=10^{-4}$ and for two different Debye lengths until a final time $T=3$, and look for the evolution of the relative energy $\mathcal{H}_{\mathcal{T}}^{n}-\mathcal{H}_{\mathcal{T}}^{\infty}$ along-time. The relative energy is decaying for all the curves plotted on Figure 6, but the velocity at which the decay occurs varies strongly depending on the Debye length and on the initial profile. In the electroneutral configurations, we mainly observe that higher quantities of $u_{0}$ yield a faster convergence. This is expected as the dissipation (14) contains a $u_{0}^{2}$ prefactor, in opposition to a mere $u_{0}$ for the classical Poisson–Nernst–Planck system. Then, still in the electroneutral regime, the smaller is $\lambda^{2}$ then the faster is the convergence towards the constant in space equilibrium.
@@
The trend is different for globally charged profiles. For the choice (91) of the initial profile, the convergence becomes extremely slow when $\lambda^{2}$ becomes small. This is due to the fact that, in some areas, the concentration of $u_{0}$ becomes extremely small. The equilibration between the other phases is then extremely slow in these regions. To illustrate this difficulty, we depict on Figure 7 the concentration profiles at $T=3$. The solvent concentration $u_{0}$ is already in good agreement with the equilibrium profile depicted on Figure 5 (see also Figure 8 for a representation of the difference between $U_{\mathcal{T},\tau}(T)$ and $U_{\mathcal{T}}^{\infty}$), but this is not the case of the other concentrations, the evolution of which being stuck in the zones where the concentration of $u_{0}$ is small. Obtaining a fine quantitative understanding of the long-time behavior of the model should be investigated in future works.